### v1

@@ Abstract
keywords:
Multimodal sarcasm analysis, large vision-language models, prompt engineering, multimodal interaction
1 Introduction
@@
2 Related Work
2.1 Multimodal Sarcasm Detection
Effective sarcasm detection plays a crucial role in tasks like sentiment analysis and opinion mining [12, 13], making it a widely researched topic. Early studies focused predominantly on sarcasm detection within the textual modality [1]. With the rise of social multimedia platforms such as Twitter and Rbbit, multimodal sarcasm detection (MSD) based on text-image pairs has gained significant attention [14]. Cai et al. [2] pioneered a hierarchical fusion-based model, integrating image, text, and image attribute features, and introduced a benchmark to evaluate its performance. Subsequently, models leveraging GNNs, OCR, object detection, and transformer-based fusion techniques have emerged [15, 16, 17, 18, 19, 20, 21, 22, 23].
Notably, Qin et al. [3] demonstrated that existing MSD models over-rely on spurious cues in text, resulting in an overestimation of their true multimodal sarcasm understanding capabilities. They introduced a more robust benchmark, MMSD2.0, for evaluating these models and proposed a Multi-view CLIP model based on CLIP’s vision-language representations to enhance sarcasm comprehension. Building on this foundation, more sophisticated models have been proposed [4, 24, 25, 26, 27]. With the rising prominence of large vision-language models (LVLMs) like LLaVA [9], Tang et al. [28] developed a demonstration-based matcher and used few-shot prompting techniques to improve MSD performance. Similarly, Wang et al. [29] introduced S3Agent, a multi-view agent framework aimed at enhancing the zero-shot MSD capabilities of LVLMs. In addition to research on text-image pairs, efforts to design models capable of handling more complex modalities (e.g., video and speech) have also emerged [30, 31, 32].
@@ 2.1 Multimodal Sarcasm Detection
2.2 Beyond Multimodal Sarcasm Detection
@@
2.3 Prompting Engineering for Sarcasm Understanding
@@ 4 Experimental Setup
Table 1: Details of selected LVLMs. In this table, we list the full names of the selected LVLMs from the HuggingFace model hub [44], along with their respective architectures. To facilitate the presentation of subsequent experimental results, we have assigned each model a short name, following the convention: {abbreviation}-{model size}.
@@
To comprehensively evaluate the performance of existing LVLMs on our proposed task, we select a diverse set of open-source models with varying architectures and scales, totaling 12 models, i.e., $N_{m}=12$. Our selection includes Idefics3 [45], InternVL 2.0 [11], Qwen2-VL [46], Phi3 [47], LLaVA-1.5 [48], LLaVA-NeXT [49], and Llama-3.2 [50]. For detailed specifications of the selected LVLMs, please refer to Table  1.
@@
5 Inter-Prompt Consistency
@@
5.1 Classification Consistency
5.1.1 Metric
As described in Section 3.1, each task yields predicted classification labels, enabling a classification-based qualitative analysis. To quantify consistency, we use Krippendorff’s $\alpha$ [52] as the Classification Consistency Score, as it effectively handles partially missing or unavailable data—particularly relevant in cases where models fail to follow instructions, resulting in missing predictions for certain samples. However, a limitation of this metric is its potential instability in scenarios with extreme class distributions or small datasets. Krippendorff’s $\alpha$ has a range of $[-1,1]$, where a value closer to 1 indicates higher consistency, and a value closer to 0 indicates poorer consistency. Negative values generally indicate that the variance in the ratings between evaluators is greater than that expected by random consistency.
5.1.2 Result
@@
5.2 Rationale Consistency
5.2.1 Metric
@@ 5.2.2 Result
6 Agreement with Ground Truth
@@ 6.3 Result
7 Model Confidence Analysis
@@ 7.2 Result
8 Neutral Label Analysis
@@ 9 Qualitative Analysis
9.1 Intra-Model Subjectivity
@@
9.2 Inter-Model Subjectivity
@@ Acknowledgments
The research presented in this paper was funded by the Anhui Provincial Natural Science Foundation under Grant No. 2308085MF220. It also received additional support from the Anhui University Natural Science Foundation through Grant Nos. 2022AH050972, 2023AH050914, and KJ2021A0516.
@@ Appendix A Detailed Results of Agreement with Ground Truth
We present the detailed results of the Agreement with Ground Truth analysis for each model across all tasks. The results for the BSC task are shown in Table 2, for the TSC task in Table 3, for the SCS task in Table 4, and for the LCS task in Table 5. Notably, as discussed in Section 6, the SCS and LCS tasks can be combined to derive the COMP results, which are illustrated in Figure 6 and further detailed in Table 6. In each table, ”voting” represents the aggregated results across all models.
Table 2: BSC results.
@@
Table 6: COMP results.
@@ Appendix B Prompt Details
B.1 Binary Sarcasm Classification
@@
B.2 Trinary Sarcasm Classification
@@
B.3 Sarcasm-Centric Scoring
@@
B.4 Literal-Centric Scoring

### v2

@@ Abstract
Index Terms:
Multimodal sarcasm analysis, large vision-language models, prompt engineering, multimodal interaction.
I Introduction
@@
II Related Work
II-A Multimodal Sarcasm Detection
Effective sarcasm detection plays a crucial role in tasks like sentiment analysis and opinion mining [12, 13, 14, 15], making it a widely researched topic. Early studies focused predominantly on sarcasm detection within the textual modality [1, 16]. With the rise of social multimedia platforms such as Twitter and Rbbit, multimodal sarcasm detection (MSD) based on text-image pairs has gained significant attention [17]. Cai et al. [2] pioneered a hierarchical fusion-based model, integrating image, text, and image attribute features, and introduced a benchmark to evaluate its performance. Subsequently, models leveraging GNNs, OCR, object detection, and transformer-based fusion techniques have emerged [18, 19, 20, 21, 22, 23, 24, 25, 26].
Notably, Qin et al. [3] demonstrated that existing MSD models over-rely on spurious cues in text, resulting in an overestimation of their true multimodal sarcasm understanding capabilities. They introduced a more robust benchmark, MMSD2.0, for evaluating these models and proposed a Multi-view CLIP model based on CLIP’s vision-language representations to enhance sarcasm comprehension. Building on this foundation, more sophisticated models have been proposed [4, 27, 28, 29, 30], leveraging enhanced multimodal interactions and efficient fine-tuning strategies for downstream tasks [31]. With the rising prominence of large vision-language models (LVLMs) like LLaVA [9], Tang et al. [32] developed a demonstration-based matcher and used few-shot prompting techniques to improve MSD performance. Similarly, Wang et al. [33] introduced S3Agent, a multi-view agent framework aimed at enhancing the zero-shot MSD capabilities of LVLMs. In addition to research on text-image pairs, efforts to design models capable of handling more complex modalities (e.g., video and speech) have also emerged [34, 35, 36].
@@ II-A Multimodal Sarcasm Detection
II-B Beyond Multimodal Sarcasm Detection
@@
II-C Prompting Engineering for Sarcasm Understanding
@@ IV Experimental Setup
TABLE I: Details of selected LVLMs.
@@
    In this table, we list the full names of the selected LVLMs from the HuggingFace model hub [54], along with their respective architectures. To facilitate the presentation of subsequent experimental results, we have assigned each model a short name, following the convention: {abbreviation}-{model size}.
@@
To comprehensively evaluate the performance of existing LVLMs on our proposed task, we select a diverse set of open-source models with varying architectures and scales, totaling 12 models, i.e., $N_{m}=12$. Our selection includes Idefics3 [48], InternVL 2.0 [11], Qwen2-VL [49], Phi3 [50], LLaVA-1.5 [51], LLaVA-NeXT [52], and Llama-3.2 [53]. For detailed specifications of the selected LVLMs, please refer to Table  I.
@@
V Inter-Prompt Consistency
@@
V-A Classification Consistency
V-A1 Metric
As described in Section III-A, each task yields predicted classification labels, enabling a classification-based qualitative analysis. To quantify consistency, we use Krippendorff’s $\alpha$ [56] as the Classification Consistency Score, as it effectively handles partially missing or unavailable data—particularly relevant in cases where models fail to follow instructions, resulting in missing predictions for certain samples. However, a limitation of this metric is its potential instability in scenarios with extreme class distributions or small datasets. Krippendorff’s $\alpha$ has a range of $[-1,1]$, where a value closer to 1 indicates higher consistency, and a value closer to 0 indicates poorer consistency. Negative values generally indicate that the variance in the ratings between evaluators is greater than that expected by random consistency.
V-A2 Result
@@
V-B Rationale Consistency
V-B1 Metric
@@ V-B2 Result
VI Agreement with Ground Truth
@@ VI-C Result
VII Model Confidence Analysis
@@ VII-B Result
VIII Neutral Label Analysis
@@ IX Qualitative Analysis
IX-A Intra-Model Subjectivity
@@
IX-B Inter-Model Subjectivity
@@ Acknowledgments
The research presented in this paper was funded by the Anhui Provincial Natural Science Foundation under Grant No. 2308085MF220. It also received additional support from the Anhui University Natural Science Foundation through Grant No. 2023AH050914.
@@ Appendix A Detailed Results of Agreement with Ground Truth
We present the detailed results of the Agreement with Ground Truth analysis for each model across all tasks. The results for the BSC task are shown in Table II, for the TSC task in Table III, for the SCS task in Table IV, and for the LCS task in Table V. Notably, as discussed in Section VI, the SCS and LCS tasks can be combined to derive the COMP results, which are illustrated in Figure 6 and further detailed in Table VI. In each table, ”voting” represents the aggregated results across all models.
TABLE II: BSC results.
@@
TABLE VI: COMP results.
@@ Appendix B Prompt Details
B-A Binary Sarcasm Classification
@@
B-B Trinary Sarcasm Classification
@@
B-C Sarcasm-Centric Scoring
@@
B-D Literal-Centric Scoring